formulate a cost function alone with constraints ,dobt need to solve it. two electric generators are interconnected to provide total power to meet the load. suppose each generatir’s cost C is a function of its power output P, and costs per unit are given by : C1=1- P1-P1^2 for generator 1. C2=1+0.6 P2+ P2^2 for generator 2. if the total power needed is at least 60 units, formulate a min-cost design problem to determine the power outputs P1 and P2 from the two units
Answer #1

Dear Student,

There is problem in the cost function of generator 1.

As cost for 1 unit power generation :-

C1= 1-1-12 = 1-1-1= -1 $

which is practically not possible as cost can not be negative.

Please recheck the question and post again.

Thank You.